# aims-PAX: Parallel Active eXploration for the automated construction of Machine Learning Force Fields

### v1

@@ II.3 Computational Details
All DFT calculations were performed using FHI aimsBlum et al. (2009) version 241114 compiled as a library and called through the python ASI package asi4pyStishenko et al. (2023) version 1.3.18 connected with ASEHjorth Larsen et al. (2017) version 3.23.0. For MACEBatatia et al. (2022, 2025) we used mace-torch version 0.3.9. with pytorchPaszke et al. (2019) version 2.3.1. For aims-PAX with PARSL, we used PARSL version 2024.12.16 and FHI aimsBlum et al. (2009) compiled as an executable. In this implementation, the ASEHjorth Larsen et al. (2017) calculator is used to perform DFT calculations.
@@
During AL, we employed the Perdew-Burke-Ernzerhof (PBE) functionalPerdew, Burke, and Ernzerhof (1996) with non-local many-body dispersion (nl-MBD)Hermann and Tkatchenko (2020) using the LIGHT species defaults for numerical settings and basis sets. Relativistic corrections were applied using the atomic ZORA approximation.van Lenthe, Snijders, and Baerends (1996) The total energy, eigenvalue, density, and force convergence criteria were set to $10^{-6}$ eV, $10^{-4}$ eV, $10^{-5}$ e/Å3, and $10^{-4}$ eV/Å, respectively. For recomputing the dataset to a higher level of theory, the PBE0Adamo, Cossi, and Barone (1999) functional with the nl-MBD dispersion correction and the INTERMEDIATE species defaults for basis sets and numerical settings was used, keeping the other settings fixed.
@@
The serial version of aims-PAX was used for both IDG and AL. The former was performed by sampling 8 points for each member of an ensemble of 4 models with a stopping criterion of a maximum of 50 epochs. The structures were sampled using the small MACE-MP0 GP modelBatatia et al. (2024) by running MD in the NVT ensemble with the Langevin thermostatBerendsen et al. (1984) at 500 K with a timestep of 1 fs and a friction coefficient of $0.001\leavevmode\nobreak\ \mathrm{fs^{-1}}$. In order to decorrelate the data points, structures were picked every 20th MD step. Their energies and forces where then computed using DFT.
@@
The AL workflow was run until a training set size of 500 structures was reached with 1:1 ratio for the validation set. During the AL, when new data was added to the training set, the models were trained for a total of 10 epochs on the updated dataset. The training was split into two steps, each involving 5 epochs. More precisely, this means that after 5 epochs are trained, the other running trajectories are propagated first before finishing with 5 more epochs of training. For more details on this see Section A.3.1. During AL, the structures were sampled using the same MD settings as in the IDG. The uncertainty threshold parameter $c_{x}$ (see Eq. 2) was set to the default value of 0.
@@
The test sets for Ac-F-A5-K were created by running MD with the small MACE-OFFKovács et al. (2025) potential in the NVT ensemble at 300, 500, and 700 K using the LangevinBerendsen et al. (1984) thermostat with a friction coefficient of $0.001\leavevmode\nobreak\ \mathrm{fs^{-1}}$ and a time step of 1 fs for 1 ns. Every 100th geometry was selected and from the remaining points 1000 were selected by farthest point samplingEldar et al. (1997) using the mean, invariant atomic MACE-OFF descriptors. The chosen geometries were then recomputed using FHI aims with the same functional and settings used in the AL.
@@
The MD simulations for assessing the stability of models were performed in the NVT ensemble at 300, 500, and 700 K using the Langevin thermostat with a friction coefficient of $0.001\leavevmode\nobreak\ \mathrm{fs^{-1}}$ and a time step of 1 fs for 1 ns. Throughout the simulation, the bond lengths were monitored, and if any of them exceeded 2 Å, the simulation was stopped.
@@
During the AL, the PBE functional with the pairwise Tkatchenko-Scheffler dispersion correctionTkatchenko and Scheffler (2009) was used, employing the INTERMEDIATE species defaults for numerical settings and basis sets. Relativistic corrections were applied using the atomic ZORA approximation.van Lenthe, Snijders, and Baerends (1996) The total energy, eigenvalue, density, and force convergence criteria were set to $10^{-6}$ eV, $10^{-5}$ eV, $10^{-5}$ e/Å3, and $10^{-4}$ eV/Å, respectively. A Gaussian smearing of $0.05$ eV was applied to the orbital occupations. The calculations employed a parallel KS method with load balancing and local indexing enabled. A maximum of $300$ self-consistency iterations was allowed. The charge mixing parameter was set to $0.02$. The k grid was set to $1\times 1\times 1$. The lattice vectors were [17.23958, 0, 0], [0, 17.23958, 0], and [0, 0, 25.00256], all in Ångstrom.
The parallel version of aims-PAX employing Parsl was used for both IDG and AL. The former was performed by sampling 10 points for each member of an ensemble of 4 models with a stopping criterion of 50 epochs for the initial training. The structures were sampled using the small MACE-MP0 GP modelBatatia et al. (2024) by running NPT MD with the Nosé-Hoover thermostatEvans and Holian (1985) at 300 K and the Parinello-Rahman barostatParrinello and Rahman (1981) at 1 bar. The timestep was set to 1 fs and otherwise default ASE parameters were used. In order to decorrelate the sampled data points, structures were picked every 20th MD step. Their energies, forces and stress where then computed using DFT. The models were converged on the initial dataset before continuing with AL by training them until no improvements w.r.t. the validation set was achieved for 50 epochs
The AL workflow was run until a training set size of 100 structures was reached with 7:3 ratio for the validation set. The maximum epochs per trajectory are 10, and the intermediate epochs are 10. The structures were sampled using the same MD settings as described in the IDG above. The uncertainty threshold parameter $c_{x}$ (see Eq. 2) was set to 0.2 and the uncertainty itself was checked every 10th MD step.
@@
System
Parameter Ac-F-A5-K (small) Ac-F-A5-K (large) CsPbI3
Channels 32 256 64
Max degree $L_{\text{max}}$ 1 2 1
Cutoff [Å] 6 6 6
Radial Bessel functions 8 8 10
Message-passing layers 2 2 2
Correlation order 3 3 3
Radial MLP layers 3 3 3
Neurons per MLP layer 32 64 16
Activation function SiLU SiLU SiLU
Output Layer Irreps "128x0e" "16x0e" "16x0e"
For the training during AL with aims-PAX the follwing settings were used. The AMSGrad optimizer Reddi, Kale, and Kumar (2018) and a learning rate of 0.01 were utilized throughout. For the the IDG, the learning rate was decreased by 0.8 using the Reduce On Plateau scheduler with a patience of 5 and $\gamma=0.9993$. No learning rate scheduler was used during the AL. An exponential moving average of 0.99 for the model parameters and a gradient clipping of 10 were used. For the loss function a weighted mean square loss of energies and forces with weights 1 and 1000, respectively, was utilized. A batch size of 5 was used throughout. After the AL runs themselves, the best performing models of the respective ensembles were trained on the final training set until there was no improvement w.r.t. the validation set for 50 epochs.
For training the large model for Ac-F-A5-K (third column in Table 1) from scratch on the recomputed dataset, the same settings as described above for aims-PAX were used except the following changes. The energy and force weights of the loss function were set to 44 and 1000, respectively. The model was trained for 1000 epochs and after the 750 epochs the energy and force weights were swapped and the learning rate set to 0.001. The learning rate was decreased by 0.8 using the Reduce On Plateau scheduler with a patience of 256 and $\gamma=0.9993$.

### v2

@@ IV Methods
DFT calculations were performed using FHI-aimsBlum et al. (2009) version 241114 compiled as a library and called through the python ASI package asi4pyStishenko et al. (2023) version 1.3.18 connected with ASEHjorth Larsen et al. (2017) version 3.26.0. For MACEBatatia et al. (2022, 2025) we used mace-torch version 0.3.9. with pytorchPaszke et al. (2019) version 2.3.1. For aims-PAX with PARSL, we used PARSL version 2024.12.16 and FHI-aimsBlum et al. (2009) compiled as an executable. In this implementation, the ASEHjorth Larsen et al. (2017) calculator is used to perform DFT calculations.
@@
During AL, we employed the Perdew-Burke-Ernzerhof (PBE) functionalPerdew, Burke, and Ernzerhof (1996) with non-local many-body dispersion (MBD-NL)Hermann and Tkatchenko (2020) using the LIGHT species defaults for numerical settings and basis sets. Relativistic corrections were applied using the atomic ZORA approximation.van Lenthe, Snijders, and Baerends (1996) The total energy, eigenvalue, density, and force convergence criteria were set to $10^{-6}$ eV, $10^{-4}$ eV, $10^{-5}$ e/Å3, and $10^{-4}$ eV/Å, respectively. For recomputing the dataset to a higher level of theory, the PBE0Adamo, Cossi, and Barone (1999) functional with the MBD-nl dispersion method and the INTERMEDIATE species defaults for basis sets and numerical settings was used, keeping the other settings fixed.
@@
The serial version of aims-PAX was used for both IDG and AL. The former was performed by sampling 8 points for each member of an ensemble of 4 models with a stopping criterion of a maximum of 50 epochs. The structures were sampled using the small MACE-MP0 GP modelBatatia et al. (2024) by running MD in the NVT ensemble with the Langevin thermostatBerendsen et al. (1984) at 500 K with a timestep of 1 fs and a friction coefficient of $0.001\penalty 10000\ \mathrm{fs^{-1}}$. In order to decorrelate the data points, structures were picked every 20th MD step. Their energies and forces were then computed using DFT.
@@
The AL workflow was run until a training set size of 500 structures was reached with a 1:1 ratio for the validation set. During the AL, when new data was added to the training set, the models were trained for a total of 10 epochs on the updated dataset. The training was split into two steps, each involving 5 epochs. More precisely, this means that after 5 epochs are trained, the other running trajectories are propagated first before finishing with 5 more epochs of training. For more details on this, see Section S5A. During AL, the structures were sampled using the same MD settings as in the IDG. The uncertainty threshold parameter $c_{x}$ (see Eq. 2) was set to the default value of 0. The uncertainty was measured every 20th MD step.
@@
The test sets for Ac-F-A5-K were created by running MD with the small MACE-OFFKovács et al. (2025) potential in the NVT ensemble at 300, 500, and 700 K using the LangevinBerendsen et al. (1984) thermostat with a friction coefficient of $0.001\penalty 10000\ \mathrm{fs^{-1}}$ and a time step of 1 fs for 1 ns. Every 100th geometry was selected, and from the remaining points, 1000 were selected by farthest point samplingEldar et al. (1997) using the mean, invariant atomic MACE-OFF descriptors. The chosen geometries were then recomputed using FHI-aims with the same functional and settings used in the AL.
@@
The MD simulations for assessing the stability of models were performed in the NVT ensemble at 300, 500, and 700 K using the Langevin thermostat with a friction coefficient of $0.001\penalty 10000\ \mathrm{fs^{-1}}$ and a time step of 1 fs for 1 ns. Throughout the simulation, the bond lengths were monitored, and if any of them exceeded 2 Å, the simulation was stopped.
MD17
During the AL, the same settings as used in the creation of MD17Chmiela et al. (2017) were used. That is, the PBE functional with the pairwise Tkatchenko-Scheffler dispersion methodTkatchenko and Scheffler (2009) was used, employing the LIGHT species defaults for numerical settings and basis sets. The default total energy, eigenvalue, density, and force convergence criteria were used. The calculations employed a parallel KS method with load balancing.
The parallel version of aims-PAX employing Parsl was used for both IDG and AL. The former was performed by sampling 10 points for each member of an ensemble of 4 models and for each distinct chemical species. After sampling, 5 training epochs were performed before continuing sampling up to a stopping criterion of 20 epochs. The structures were sampled using the small MACE-MP0 GP modelBatatia et al. (2024) at 500 K with a timestep of 1 fs and a friction coefficient of $0.001\penalty 10000\ \mathrm{fs^{-1}}$. In order to decorrelate the data points, structures were picked every 25th MD step. Their energies and forces were then computed using DFT.
The AL workflow was run until a training set size of 1000 structures was reached with a 10:1 ratio for the validation set. During the AL, when new data was added to the training set, the models were trained for a total of 1 epoch on the updated dataset. During AL, the structures were sampled using the same MD settings as in the IDG. The uncertainty threshold parameter $c_{x}$ (see Eq. 2) was set to the default value of 0. The threshold was frozen after the training set size reached a size of 500 geometries. The uncertainty was measured every 25th MD step.
The data for training the MLFF from scratch on MD17 was obtained by taking the first 50k structures for each molecule in the original dataset. Then every 25th structure was selected from this subset, resulting in 2,000 points per species. From these 2,000 points, 100 points for training and 10 points for validation were randomly chosen per species. All of these subsets were then combined, resulting in 1,000 training and 100 validation points.
System
Parameter Ac-F-A5-K (small) / MD17 Ac-F-A5-K (large) CsPbI3 Paracetamol+H2O
Channels 32 256 64 128
Max degree $L_{\text{max}}$ 1 2 1 1
Cutoff [Å] 6 6 6 6
Radial Bessel functions 8 8 10 8
Message-passing layers 2 2 2 2
Correlation order 3 3 3 3
Radial MLP layers 3 3 3 3
Neurons per MLP layer 32 64 16 64
Activation function SiLU SiLU SiLU SiLU
Output Layer Irreps "128x0e" "16x0e" "16x0e" "128x0e"
Paracetamol in Water
During the AL, the PBE functional with many-body dispersion (MBD)Tkatchenko and Scheffler (2009) using the LIGHT species defaults for numerical settings and basis sets was used. The default total energy, eigenvalue, density, and force convergence criteria were used. The calculations employed a parallel KS method with load balancing. For the periodic system a k grid of $2\times 2\times 2$ was utilized.
The initial structures for paracetamol surrounded by a cluster of 90 water molecules, bulk water with 64 and 128 molecules at a density of 1 g/mL and paracetamol surrounded by 600 water molecules at a density of 1 g/mL were created using PackmolMartínez et al. (2009) version 21.1.0.
Paracetamol in the gas phase, bulk water with 64 molecules, and paracetamol surrounded by a water cluster were optimized using FHI-aims with a convergence threshold on the force of 0.01 eV/Å before using them in aims-PAX. Before MD production runs of paracetamol in the gas phase, bulk water with 128 molecules, and paracetamol in explicit water (600 water molecules), the structures were first optimized using FHI-aims for the former and the medium MACE-MP0 GP modelBatatia et al. (2024) for the latter two with a convergence threshold on the force of 0.01 eV/Å. Then optimization was repeated using the MACE model acquired from the aims-PAX run with a convergence threshold on the force of 0.01 eV/Å. All optimizations were performed using the Broyden–Fletcher–Goldfarb–Shanno algorithmFletcher (2000) as implemented in FHI-aims and ASEHjorth Larsen et al. (2017), respectively.
The parallel version of aims-PAX employing Parsl was used for both IDG and AL. The former was performed by sampling 5 points for each member of an ensemble of 4 models and for each trajectory. After sampling, 5 training epochs were performed before continuing sampling up to a stopping criterion of 20 epochs. The structures were sampled using the medium MACE-MP0 GP modelBatatia et al. (2024) from 7 trajectories in total. Of these, one was of paracetamol in the gas phase at 300 K (NVT) with a timestep of 1 fs and a friction coefficient of 0.001 fs-1; three were of paracetamol surrounded by a cluster of 90 water molecules at 300 K, 350 K, and 400 K (NVT) with a timestep of 0.5 fs and a friction coefficient of 0.001 fs-1; and the remaining three were of bulk water at 1 atm (NPT) and 300 K, 400 K, and 500 K with a timestep of 0.5 fs using full Martyna-Tobias-Klein (MTK) dynamicsMartyna, Tobias, and Klein (1994) as implemented in ASEHjorth Larsen et al. (2017) with a temperature and pressure damping factor of 100 and 1000, respectively. In order to decorrelate the data points, structures were picked every 25th MD step. Their energies and forces were then computed using DFT.
The AL workflow was run until a training set size of 1,000 structures was reached with a 10:1 ratio for the validation set. During the AL, when new data was added to the training set, the models were trained for a total of 1 epoch on the updated dataset. During AL, the structures were sampled using the same MD settings as in the IDG. The uncertainty threshold parameter $c_{x}$ (see Eq. 2) was set to the default value of 0 and the uncertainty was measured every 50th MD step. The threshold was frozen after the training set size reached a size of 500 geometries.
In order to acquire the vibrational density of states (VDoS) of paracetamol in the gas phase, the following protocol was applied. First, 20 starting geometries were picked through k-means clustering from the MD17 dataset using the dihedral angles $\tau_{1},\tau_{2},\tau_{3}$ (see Fig. 4b) as the descriptor. From the 20 clusters, the geometry closest to the respective cluster centers was used as a starting point for a 10 ps NVT simulation at 300 K with a timestep of 1 fs and a friction coefficient of 0.001 fs-1. The final structures and their velocities were then used to perform MD runs in the NVE ensemble with a timestep of 0.1 fs. Subsequently, the velocity autocorrelation function (ACF) was computed from the combined trajectories, and the VDoS was acquired through a Fourier transform of the velocity ACF.
The oxygen-oxygen radial distribution function (RDF) of bulk water was acquired from a 500 ps MD simulation of 128 water molecules with periodic boundary conditions, whereas the first 200 ps were used to equilibrate the system. The simulation was performed in the NPT ensemble using MTK dynamics at 330K and 1 atm. The RDF was computed using the analysis tools implemented in ASEHjorth Larsen et al. (2017).
In order to acquire the distributions of dihedral angles $\tau_{1},\tau_{2},\tau_{3}$ (see Fig. 4b) of paracetamol in the gas phase, 20 independent 800 ps long MD simulations in the NVT ensemble from the optimized geometry at 300 K with a timestep of 0.5 fs and a friction coefficient of 0.001 fs-1 were performed. The first 50 ps were used for equilibration. In case of paracetamol in water, 36 independent 800 ps long MD simulations in the NVT ensemble at 300 K with a timestep of 0.5 fs and a friction coefficient of 0.001 fs-1 were performed. The starting geometries were chosen from a separate MD run so that the dihedral angles $\tau_{2}$ and $\tau_{3}$ were equally distributed. The first 50 ps were used for equilibration.
@@
During the AL, the PBE functional with the pairwise Tkatchenko-Scheffler dispersion methodTkatchenko and Scheffler (2009) was used, employing the INTERMEDIATE species defaults for numerical settings and basis sets. Relativistic corrections were applied using the atomic ZORA approximation.van Lenthe, Snijders, and Baerends (1996) The total energy, eigenvalue, density, and force convergence criteria were set to $10^{-6}$ eV, $10^{-5}$ eV, $10^{-5}$ e/Å3, and $10^{-4}$ eV/Å, respectively. A Gaussian smearing of $0.05$ eV was applied to the orbital occupations. The calculations employed a parallel KS method with load balancing and local indexing enabled. A maximum of $300$ self-consistency iterations was allowed. The charge mixing parameter was set to $0.02$. The k grid was set to $1\times 1\times 1$. The lattice vectors were [17.23958, 0, 0], [0, 17.23958, 0], and [0, 0, 25.00256], all in Ångstrom.
The parallel version of aims-PAX employing Parsl was used for both IDG and AL. The former was performed by sampling 10 points for each member of an ensemble of 4 models with a stopping criterion of 50 epochs for the initial training. The structures were sampled using the small MACE-MP0 GP modelBatatia et al. (2024) by running NPT MD with the Nosé-Hoover thermostatEvans and Holian (1985) at 300 K and the Parinello-Rahman barostatParrinello and Rahman (1981) at 1 bar. The timestep was set to 1 fs and otherwise default ASEHjorth Larsen et al. (2017) parameters were used. In order to decorrelate the sampled data points, structures were picked every 20th MD step. Their energies, forces, and stress were then computed using DFT. The models were converged on the initial dataset before continuing with AL by training them until no improvements w.r.t. the validation set were achieved for 50 epochs
The AL workflow was run until a training set size of 100 structures was reached with a 7:3 ratio for the validation set. The maximum epochs per trajectory were 10, and the intermediate epochs were 10. The structures were sampled using the same MD settings as described in the IDG above. The uncertainty threshold parameter $c_{x}$ (see Eq. 2) was set to 0.2, and the uncertainty itself was checked every 10th MD step.
@@
For the training during AL with aims-PAX the following settings were used. The AMSGrad optimizer Reddi, Kale, and Kumar (2018) and a learning rate of 0.01 were utilized throughout. For the IDG, the learning rate was decreased by 0.8 using the Reduce On Plateau scheduler with a patience of 5 and $\gamma=0.9993$. No learning rate scheduler was used during the AL. An exponential moving average of 0.99 for the model parameters and a gradient clipping of 10 were used. For the loss function, a weighted mean square loss of energies and forces with weights 1 and 1000, respectively, was utilized. A batch size of 5 was used throughout. After the AL runs themselves, the best-performing models of the respective ensembles were trained on the final training set until there was no improvement w.r.t. the validation set for 50 epochs.
For training the large model for Ac-F-A5-K (third column in Table 1) from scratch on the recomputed dataset, the same settings as described above for aims-PAX were used except for the following changes. The energy and force weights of the loss function were set to 44 and 1000, respectively. The model was trained for 1000 epochs, and after 750 epochs, the energy and force weights were swapped and the learning rate set to 0.001. The learning rate was decreased by 0.8 using the Reduce On Plateau scheduler with a patience of 256 and $\gamma=0.9993$.
For training the model for MD17 from scratch the same settings as described above for aims-PAX were used except for the following changes. The model was trained for a total of 500 epochs. The MACE architecture was the one summarized in the first column in Table 1.